# On dense orbits in the space of subequivalence relations

### v2

@@ Abstract
We first explain how to endow the space of subequivalence relations of any non-singular countable equivalence relation with a Polish topology, extending the framework of Kechris’ recent monograph on subequivalence relations of probability measure-preserving (p.m.p.) countable equivalence relations. We then restrict to p.m.p. equivalence relations and discuss dense orbits therein for the natural action of the full group and of the automorphism group of the relation. Our main result is a characterization of the subequivalence relations having a dense orbit in the space of subequivalence relations of the ergodic hyperfinite p.m.p. equivalence relation. We also show that in this setup, all full groups orbits are meager. We finally provide a few Borel complexity calculations of natural subsets in spaces of subequivalence relations using a natural metric we call the uniform metric. This answers some questions from an earlier version of Kechris’ monograph.
@@ 1 Introduction
[…]
Our next result is in the p.m.p. setup, and was motivated by the question of dense orbits in the space of subequivalence relations, asked by Kechris in his monograph. Indeed, it is very natural to try to understand subequivalence relations up to isomorphism, and this is precisely what the action of the automorphism group of $\mathcal{R}$ on subequivalence relations of $\mathcal{R}$ encodes. As it turns out, when $\mathcal{R}$ is the ergodic hyperfinite p.m.p. equivalence relation we obtain the following complete characterization of dense orbits for the natural actions of both the full group $[\mathcal{R}]$ and the automorphism group $\mathrm{Aut}(\mathcal{R})$.
@@
    The $[\mathcal{R}_{0}]$-orbit of $\mathcal{S}$ is dense in $\operatorname{\mathrm{Sub}}(\mathcal{R}_{0})$;
@@
    the closure of the $[\mathcal{R}]$-orbit of $\mathcal{S}$ contains $\Delta_{X}$;
@@
Let $\mathcal{R}_{0}$ be the ergodic p.m.p. hyperfinite equivalence relation. Then all $[\mathcal{R}_{0}]$-orbits are meager in $\operatorname{\mathrm{Sub}}(\mathcal{R}_{0})$.
It then follows from Mycielcki’s theorem (see e.g. [Gao09, Thm. 5.3.1]) that $\mathcal{R}_{0}$ contains a continuum of aperiodic subequivalence relations of infinite index which are in pairwise disjoint full group orbits. Moreover, by [Kec, Prop. 10.1] and Example 1.2, such a continuum can be found inside ergodic subequivalence relations. We leave the following question open.
@@
This question can also be asked for equivalence relations $\mathcal{R}$ different from $\mathcal{R}_{0}$, although the existence of dense $\mathrm{Aut}(\mathcal{R})$- or $[\mathcal{R}]$-orbits in their space of subequivalence relations is already an open problem (see the remark preceding Corollary 4.13 for some examples where there are no dense orbits though). Also note that if $\mathcal{R}$ is a type II or type III ergodic non-singular equivalence relation, it is not hard to show that there are always dense full group orbits in $\operatorname{\mathrm{Sub}}(\mathcal{R})$ (see the remark right before Section 4.4 for a sketch of proof), but the question of comeager orbits is completely open there.
@@
[…]
@@ Acknowledgements.
I would like to thank Alekos Kechris for his remarks and encouragements on this paper, and Todor Tsankov for discussions around this topic. I am very much indebted to my coauthors Pierre Fima, Kunal Mukherjee and Issan Patri for many of the ideas in this paper, which were developped in the parallel framework of subalgebras of finite von Neumann algebras. I am also thankful to Damien Gaboriau for his comments, and to Stefaan Vaes for allowing me to include his remark on the von Neumann algebraic way of understanding the space of subequivalence relations.
@@ 2.1 The measure algebra of a standard $\sigma$-finite space
Let $(A_{n})$ be a sequence of elements of $\mathrm{MAlg}(Y,\mu)$ such that $\sum_{n}\mu(A_{n}\bigtriangleup A_{n+1})<+\infty$. Let lim $\liminf A_{n}=\bigcup_{N}\bigcap_{n\geqslant N}A_{n}$, then the sequence $(A_{n})$ converges to lim $\liminf A_{n}$.
@@ 2.2 The space of subequivalence relations
Given an action of a countable group $\Gamma$ by Borel bijections on $X$ standard Borel, we obtain the associated equivalence relation $\mathcal{R}_{\Gamma}=\{(x,\gamma x)\colon x\in X,\gamma\in\Gamma\}$ whose equivalence classes are exactly the $\Gamma$-orbits. This equivalence relation is a Borel subset of $X\times X$, and its classes are countable: $\mathcal{R}_{\Gamma}$ is by definition a countable Borel equivalence relation (CBER, which can conveniently be read as “seeber"). Conversely, the Feldman-Moore theorem ensures us that every CBER comes from a Borel action of a countable group, see for instance [KM04, Thm. 1.3].
Given such an equivalence relation, define its Borel pseudo full group as the set of all partial bijection $\varphi:\mathrm{dom}\;\varphi\subseteq X\to\operatorname{\mathrm{rng}}\varphi\subseteq X$ such that for all $x\in A$, we have $(x,\varphi(x))\in\mathcal{R}$. By definition, the Borel full group is the group made of elements of the Borel pseudo full group whose domain and range is equal to $X$.
@@
A Borel subset of $\mathcal{R}$ is called a Borel subequivalence relation if it defines an equivalence relation on $X$. An element of $\mathrm{MAlg}(\mathcal{R},M)$ is called a subequivalence relation if its equivalence class contains an equivalence relation. Denote by $\operatorname{\mathrm{Sub}}(\mathcal{R})\subseteq\mathrm{MAlg}(\mathcal{R},M)$ the space of subequivalence relations of $\mathcal{R}$. We can now easily prove the following result of Kechris (our proof is moreover motivated by [Kec, Thm. 5.1], which we have essentially recast as Proposition 2.3).
@@
[…]
@@
By definition for almost all $x$ we have $[x]_{\mathcal{S}_{1}}=[x]_{\mathcal{S}_{2}}$. Since $\mathcal{S}_{1}$ is non singular we find a smaller $\mathcal{S}_{1}$-invariant set $X_{0}$ for which the same conclusion holds (if $\mathcal{S}_{1}=\mathcal{R}_{\Gamma}$ we let $X_{0}=\bigcap_{\gamma\in\Gamma}\gamma X_{1}$, where $X_{1}=\{x\in X\colon[x]_{\mathcal{S}_{1}}=[x]_{\mathcal{S}_{2}}\}$). This finishes the proof. ∎
@@
Given a non-singular equivalence relation $\mathcal{R}$ on $(X,\mu)$, consider the space of subalgebras of its von Neumann algebra $L\mathcal{R}$, which is Polish for the Maréchal topology ([Mar73], see also [FMMP24] for a detailed proof). Then it is a result of Aoi that subequivalence relations of $\mathcal{R}$ are in one-to-one correspondence with subalgebras of $L\mathcal{R}$ containing $\mathrm{L}^{\infty}(X,\mu)$ [Aoi03]. It is not hard to see that such subalgebras form a closed set, and the map which associates to $\mathcal{S}\in\operatorname{\mathrm{Sub}}(\mathcal{R})$ the subalgebra $L\mathcal{S}$ of $L\mathcal{R}$ is a homeomorphism (see Theorem E from the recent preprint by Shuoxing Zhou [Zho24]), yielding yet another way of understanding the Polish topology induced by the measure algebra of $\mathcal{R}$ on $\operatorname{\mathrm{Sub}}(\mathcal{R})$. For (many!) other interpretations of this topology in the probability measure preserving setup, the reader should consult [Kec].
@@ 3 Strong and lower topologies
In this section, we first study the strong topology on the space of subequivalence relations as defined in [Kec], which will be useful towards establishing our orbit density result (Theorem B). We then use a lower topology on the measure algebra so as to obtain a natural proof of the fact that the space of subequivalence relations is closed in the measure algebra, and of the fact that the map which takes a graphing to the equivalence it generates is Baire class one [Kec, Prop. 19.1].
@@ 3.1 Topologies on measure algebras of $\sigma$-finite spaces
Observe that every element $\varphi$ of the Borel pseudo full group of a CBER $\mathcal{R}$ is completely determined by its graph, namely the set $\{(x,\varphi(x))\}$, which is a Borel subset of $\mathcal{R}$. From now on, we identify elements of the Borel pseudo full group to their graph. Doing this up to measure zero when $\mathcal{R}$ is nonsingular, we arrive at the following definition.
@@
The pseudo full group of a non-singular equivalence relation $\mathcal{R}$ on $(X,\mu)$ is the quotient of the Borel pseudo full group by the equivalence relation which identifies $\varphi_{1}$, $\varphi_{2}$ when $M(\varphi_{1}\bigtriangleup\varphi_{2})=0$. It is denoted by $[[\mathcal{R}]]$, and for ease of notation we still write its elements as $\varphi$’s. Finally the full group of $\mathcal{R}$ denotes all the elements of $[[\mathcal{R}]]$ with full measure domain and range.
[…]
We now make a modification of Kechris’ uniquely generating sequences of involutions. Let us say that an element $\varphi$ of the Borel pseudo-full group is a moving partial involution if $\mathrm{dom}\;\varphi=\operatorname{\mathrm{rng}}\varphi$ and for all $x\in\mathrm{dom}\;\varphi$, we both have $\varphi(x)\neq x$ and $\varphi(\varphi(x))=x$.
@@
A sequence $(\varphi_{n})$ of moving partial involutions of the Borel pseudo-full group of a CBER $\mathcal{R}$ is called a uniquely generating sequence of moving partial involutions if
$\mathcal{R}\setminus\Delta_{X}=\bigsqcup_{n}\varphi_{n},$
@@
Let $(T_{n})$ be a uniquely generating sequence of involutions for $\mathcal{R}$ as in [Kec, Prop. 4.6], and let $\varphi_{n}$ be the restriction of $T_{n}$ to its support. ∎
@@
$\operatorname{\mathrm{Sub}}(\mathcal{R})\to\prod_{n\in\mathbb{N}}\mathrm{MAlg}(\varphi_{n},M_{\varphi_{n}}))$
which takes $\mathcal{S}$ to the sequence $(\mathcal{S}\cap\varphi_{n})_{n}$ is a homeomorphism onto its image.
@@ 3.2 Lower topologies on measure algebras
Let us first work a bit more on the relationship between $M$ and $m$. Observe that $M$ defines a lower-semi continuous function on $\mathrm{MAlg}(Y,M)$ (possibly taking +$\infty$+ ∞ as a value). Indeed if $M(A)>\alpha$, we have a $\delta>0$ such that $M(C)<M(A)-\alpha$ whenever $C$ is a subset of $A$ satisfying $m(C)<\delta$, and so as soon as $m(A\setminus B)<\delta$ we have $M(B)>\alpha$. In particular for every $k\in\mathbb{N}$ the set of elements with measure $\leqslant k$ is closed.
@@
Going back to the case $Y=\mathcal{R}$, observe that if $T$ in the full group of $\mathcal{R}$ is aperiodic, then $T^{n}\to\emptyset$ for $d_{m}$ while it stays in the sphere around $1$ for $d_{M}$. In particular $d_{m}$ does not refine $d_{M}$ (here again we identified an element of the full group to its graph). Let us also mention that $[\mathcal{R}]$ is $d_{M}$-closed, and that the metric induced by $d_{M}$ on $[\mathcal{R}]$ is the well-known uniform metric $d_{u}(S,T)=\mu(\{x\in X\colon S(x)\neq T(x)\})$, endowing $[\mathcal{R}]$ with a Polish group topology.
@@
Let $(Y,M)$ be a $\sigma$-finite measured space (the measure $M$ can be finite). We endow its measure algebra with the lower $M$-topology which is defined by taking as neighborhoods of an element $A\in\mathrm{MAlg}(Y,M)$ all sets of the form
@@
[…]
@@
Let $\epsilon>0$. Note that $(A_{1}\cap A_{2})\setminus(B_{1}\cap B_{2})\subseteq(A_{1}\cap A_{2})\setminus B_{1})\cup((A_{1}\cap A_{2})\setminus B_{2})\subseteq(A_{1}\setminus B_{1})\cup(A_{2}\setminus B_{2})$, so if if $M(A_{1}\setminus B_{1})<\epsilon$ and $M(A_{2}\setminus B_{2})<\epsilon$, then $M((A_{1}\cap A_{2})\setminus(B_{1}\cap B_{2}))<2\epsilon$ as wanted. ∎
@@
Let $(A_{n})_{n\in\mathbb{N}}$ be a family of elements of the measure algebra, then since $m$ is finite we may find $N$ such that $\mu((\bigcup_{n}A_{n})\setminus\bigcup_{n<N}A_{n}))<\epsilon$. Now if we take $B_{0},...,B_{N-1}$ such that $m(A_{i}\setminus B_{i})<\epsilon/N$, and any $B_{N},B_{N+1},...$, we see that $\mu((\bigcup_{n}A_{n})\setminus(\bigcup_{n}B_{n}))<2\epsilon$ as wanted. ∎
@@
Note that the countable union map is not continuous if we put the usual topology everywhere (indeed $X$ can easily be obtained as a limit of reunions of sequences of sets which converge coordinatewise to $\emptyset$). Here in nevertheless a positive result for the usual topology which will be useful later on.

### v3

@@ Abstract
We first explain how to endow the space of subequivalence relations of any non-singular countable equivalence relation with a Polish topology, extending the framework of Kechris’ recent monograph on subequivalence relations of probability measure-preserving (p.m.p.) countable equivalence relations. We then restrict to p.m.p. equivalence relations and discuss dense orbits therein for the natural action of the full group and of the automorphism group of the relation. Our main result is a characterization of the subequivalence relations having a dense orbit in the space of subequivalence relations of the ergodic hyperfinite p.m.p. equivalence relation. We also show that in this setup, all orbits under the full group action are meager. We finally provide a few Borel complexity calculations of natural subsets in spaces of subequivalence relations using a natural metric we call the uniform metric. This answers some questions from an earlier version of Kechris’ monograph.
@@ 1 Introduction
We actually provide two proofs of Theorem A. The first one is very short and uses the fact that in a measure algebra, converging sequences always admit subsequences whose limit is actually equal to their $\liminf$ (see Proposition 2.3). This observation provides a unifying viewpoint to some special cases noted by Kechris, see [Kec, Thm. 5.1 and Thm. 18.3]. The second proof is somehow more natural in that we directly show that the various axioms of being a subequivalence relation define closed subsets in the measure algebra. Along the way, we study a natural non-Hausdorff topology on the measure algebra that we call the lower topology. This topology has some nice continuity properties which also allow us to give a slick proof of the fact that the map which associates to a subset of an equivalence relation the equivalence relation it generates is Baire class one (see Proposition 3.19, which is a natural generalization of [Kec, Prop. 19.1]).
Our next result is in the p.m.p. setup, and was motivated by the question of dense orbits in the space of subequivalence relations, asked by Kechris in his monograph. Indeed, it is very natural to try to understand subequivalence relations up to isomorphism, and this is precisely what the action of the automorphism group of $\mathcal{R}$ on subequivalence relations of $\mathcal{R}$ encodes. As it turns out, when $\mathcal{R}$ is the ergodic hyperfinite p.m.p. equivalence relation we obtain the following complete characterization of dense orbits for the natural actions of both the full group $\mathcal{R}$] and the automorphism group $\mathrm{Aut}(\mathcal{R})$.
@@
    The $\mathcal{R}_{0}]-$orbit of $\mathcal{S}$ is dense in $\operatorname{\mathrm{Sub}}(\mathcal{R}_{0})$;
@@
    the closure of the $\mathcal{R}$]-orbit of $\mathcal{S}$ contains $\Delta_{X}$;
@@
Let $\mathcal{R}_{0}$ be the ergodic p.m.p. hyperfinite equivalence relation. Then all $\mathcal{R}_{0}]-$orbits are meager in $\operatorname{\mathrm{Sub}}(\mathcal{R}_{0})$.
It then follows from Mycielski’s theorem (see e.g. [Gao09, Thm. 5.3.1]) that $\mathcal{R}_{0}$ contains a continuum of aperiodic subequivalence relations of infinite index which are in pairwise disjoint full group orbits. Moreover, by [Kec, Prop. 10.1] and Example 1.2, such a continuum can be found inside ergodic subequivalence relations. We leave the following question open.
@@
This question can also be asked for equivalence relations $\mathcal{R}$ different from $\mathcal{R}_{0}$, although the existence of dense $\mathrm{Aut}(\mathcal{R})$- or $\mathcal{R}$]-orbits in their space of subequivalence relations is already an open problem (see the remark preceding Corollary 4.13 for some examples where there are no dense orbits though). Also note that if $\mathcal{R}$ is a type II or type III ergodic non-singular equivalence relation, it is not hard to show that there are always dense full group orbits in $\operatorname{\mathrm{Sub}}(\mathcal{R})$ (see the remark right before Section 4.4 for a sketch of proof), but the question of comeager orbits is completely open there.
@@
Let us finish by outlining the paper: along with basic preliminaries, Section 2 provides a first proof of Theorem A using the $\liminf$. Section 3 contains the second proof, which relies in a more direct way on the topology of the measure algebra, using as well the lower topology that we define and study therein. Section 4 is devoted to the proof of Theorem B and D in their more general form, and contains in particular the proof of Theorem C in Subsection 4.2. Finally, Section 5 contains our complexity computations, which use the uniform metric.
@@ Acknowledgements.
I would like to thank Alekos Kechris for his remarks and encouragements on this paper, and Todor Tsankov for discussions around this topic. I am very much indebted to my coauthors Pierre Fima, Kunal Mukherjee and Issan Patri for many of the ideas in this paper, which were developped in the parallel framework of subalgebras of finite von Neumann algebras. I am also thankful to Damien Gaboriau for his comments, and to Stefaan Vaes for allowing me to include his remark on the von Neumann algebraic way of understanding the space of subequivalence relations. Finally, I am very grateful to the anonymous referee for their helpful comments which improved the paper in many ways.
@@ 2.1 The measure algebra of a standard $\sigma$-finite space
Let $(A_{n})$ be a sequence of elements of $\mathrm{MAlg}(Y,\mu)$ such that $\sum_{n}\mu(A_{n}\bigtriangleup A_{n+1})<+\infty$. Let $\liminf A_{n}=\bigcup_{N}\bigcap_{n\geqslant N}A_{n}$, then the sequence $(A_{n})$ converges to $\liminf A_{n}$.
@@ 2.2 The space of subequivalence relations
Given an action of a countable group $\Gamma$ by Borel bijections on a standard Borel space $X$, we obtain the associated equivalence relation $\mathcal{R}_{\Gamma}=\{(x,\gamma x)\colon x\in X,\gamma\in\Gamma\}$ whose equivalence classes are exactly the $\Gamma$-orbits. This equivalence relation is a Borel subset of $X\times X$, and its classes are countable: $\mathcal{R}_{\Gamma}$ is by definition a countable Borel equivalence relation (CBER, which can conveniently be read as “seeber"). Conversely, the Feldman-Moore theorem ensures us that every CBER comes from a Borel action of a countable group, see for instance [KM04, Thm. 1.3].
Given such an equivalence relation, define its Borel pseudo full group as the set of all partial Borel bijection $\varphi:\mathrm{dom}\;\varphi\subseteq X\to\operatorname{\mathrm{rng}}\varphi\subseteq X$ such that for all $x\in A$, we have $(x,\varphi(x))\in\mathcal{R}$. By definition, the Borel full group is the group made of elements of the Borel pseudo full group whose domain and range is equal to $X$.
@@
A Borel subset of $\mathcal{R}$ is called a Borel subequivalence relation if it is an equivalence relation on $X$. An element of $\mathrm{MAlg}(\mathcal{R},M)$ is called a subequivalence relation if its equivalence class contains an equivalence relation. Denote by $\operatorname{\mathrm{Sub}}(\mathcal{R})\subseteq\mathrm{MAlg}(\mathcal{R},M)$ the space of subequivalence relations of $\mathcal{R}$. We can now easily prove the following result of Kechris (our proof is moreover motivated by [Kec, Thm. 5.1], which we have essentially recast as Proposition 2.3).
@@
Observe that the $\liminf$ of any sequence of equivalence relations is an equivalence relation. By Proposition 2.3 the result follows. ∎
@@
By definition for almost all $x$ we have $x]_{\mathcal{S}_{1}}=[x]_{\mathcal{S}_{2}}$. Since $\mathcal{S}_{1}$ is non singular we find a smaller $\mathcal{S}_{1}$-invariant set $X_{0}$ for which the same conclusion holds (if $\mathcal{S}_{1}=\mathcal{R}_{\Gamma}$ we let $X_{0}=\bigcap_{\gamma\in\Gamma}\gamma X_{1}$, where $X_{1}=\{x\in X\colon[x]_{\mathcal{S}_{1}}=[x]_{\mathcal{S}_{2}}\}$). This finishes the proof. ∎
@@
Given a non-singular equivalence relation $\mathcal{R}$ on $(X,\mu)$, consider the space of subalgebras of its von Neumann algebra $L\mathcal{R}$, which is Polish when endowed with the Maréchal topology ([Mar73], see also [FMMP24] for a detailed proof). Then it is a result of Aoi that subequivalence relations of $\mathcal{R}$ are in one-to-one correspondence with subalgebras of $L\mathcal{R}$ containing $\mathrm{L}^{\infty}(X,\mu)$ [Aoi03]. It is not hard to see that such subalgebras form a closed set, and the map which associates to $\mathcal{S}\in\operatorname{\mathrm{Sub}}(\mathcal{R})$ the subalgebra $L\mathcal{S}$ of $L\mathcal{R}$ is a homeomorphism (see Theorem E from the recent preprint by Shuoxing Zhou [Zho24]), yielding yet another way of understanding the Polish topology induced by the measure algebra of $\mathcal{R}$ on $\operatorname{\mathrm{Sub}}(\mathcal{R})$. For (many!) other interpretations of this topology in the probability measure preserving setup, the reader should consult [Kec].
@@ 3 Strong and lower topologies
In this section, we first study the strong topology on the space of subequivalence relations as defined in [Kec], which will be useful towards establishing our orbit density result (Theorem B). We then use a lower topology on the measure algebra so as to obtain a natural proof of the fact that the space of subequivalence relations is closed in the measure algebra, and of the fact that the map which takes a subgraph to the equivalence it generates is Baire class one [Kec, Prop. 19.1].
@@ 3.1 Topologies on measure algebras of $\sigma$-finite spaces
Observe that every element $\varphi$ of the Borel pseudo full group of a CBER $\mathcal{R}$ is completely determined by its graph, namely the set $\{(x,\varphi(x))\colon x\in\mathrm{dom}\;\varphi\}$, which is a Borel subset of $\mathcal{R}$. From now on, we identify elements of the Borel pseudo full group to their graph. Doing this up to measure zero when $\mathcal{R}$ is nonsingular, we arrive at the following definition.
@@
The pseudo full group of a non-singular equivalence relation $\mathcal{R}$ on $(X,\mu)$ is the quotient of the Borel pseudo full group by the equivalence relation which identifies $\varphi_{1}$, $\varphi_{2}$ when $M(\varphi_{1}\bigtriangleup\varphi_{2})=0$. It is denoted by $\mathcal{R}$]], and for ease of notation we still write its elements as $\varphi$’s. Finally the full group of $\mathcal{R}$ denotes all the elements of $\mathcal{R}$]] with full measure domain and range.
It is a direct consequence of the definition of $M$ that $M(\varphi_{1}\bigtriangleup\varphi_{2})=0$ iff $\mu(\mathrm{dom}\;\varphi_{1}\bigtriangleup\mathrm{dom}\;\varphi_{2})=0$ and for almost all $x\in\mathrm{dom}\;\varphi_{1}\cap\mathrm{dom}\;\varphi_{2}$, we have $\varphi_{1}(x)=\varphi_{2}(x)$. The Borel pseudo full group is stable under countable increasing unions and arbitrary countable intersections, so it is stable under taking $\liminf$. Following the same approach as for the proof of Theorem 2.6, one obtains that $\mathcal{R}$]] is a closed subset of $\mathrm{MAlg}(\mathcal{R},M)$, in particular it is Polish. Note however that $\mathcal{R}$] is not closed for this topology (see Example 3.6).
We now make a slight modification of Kechris’ uniquely generating sequences of involutions. Let us say that an element $\varphi$ of the Borel pseudo-full group is a moving partial involution if $\mathrm{dom}\;\varphi=\operatorname{\mathrm{rng}}\varphi$ and for all $x\in\mathrm{dom}\;\varphi$, we both have $\varphi(x)\neq x$ and $\varphi(\varphi(x))=x$.
@@
A sequence $(\varphi_{n})_{n\in\mathbb{N}}$ of moving partial involutions of the Borel pseudo-full group of a CBER $\mathcal{R}$ is called a uniquely generating sequence of moving partial involutions if
$\mathcal{R}\setminus\Delta_{X}=\bigsqcup_{n\in\mathbb{N}}\varphi_{n},$
@@
Let $(T_{n})_{n\in\mathbb{N}}$ be a uniquely generating sequence of involutions for $\mathcal{R}$ as in [Kec, Prop. 4.6], and for every $n\in\mathbb{N}$ let $\varphi_{n}$ be the restriction of $T_{n}$ to its support. ∎
@@
$\operatorname{\mathrm{Sub}}(\mathcal{R})\to\prod_{n\in\mathbb{N}}\mathrm{MAlg}(\varphi_{n},M_{\varphi_{n}})$
which takes $\mathcal{S}$ to the sequence $(\mathcal{S}\cap\varphi_{n})_{n\in\mathbb{N}}$ is a homeomorphism onto its image.
@@ 3.2 Lower topologies on measure algebras
Let us first work a bit more on the relationship between $M$ and $m$. Observe that $M$ defines a lower-semi continuous function on $\mathrm{MAlg}(Y,M)$ (possibly taking $\infty$ as a value). Indeed if $M(A)>\alpha$, we have a $\delta>0$ such that $M(C)<M(A)-\alpha$ whenever $C$ is a subset of $A$ satisfying $m(C)<\delta$, and so as soon as $m(A\setminus B)<\delta$ we have $M(B)>\alpha$. In particular for every $k\in\mathbb{N}$ the set of elements with measure $\leqslant k$ is closed.
@@
Going back to the case $Y=\mathcal{R}$, observe that if $T$ in the full group of $\mathcal{R}$ is aperiodic, then $T^{n}\to\emptyset$ for $d_{m}$ while it stays in the sphere around $1$ for $d_{M}$. In particular $d_{m}$ does not refine $d_{M}$. Let us also mention that $\mathcal{R}$] is $d_{M}$-closed, and that the metric induced by $d_{M}$ on $\mathcal{R}$] is the well-known uniform metric $d_{u}(S,T)=\mu(\{x\in X\colon S(x)\neq T(x)\})$, endowing $\mathcal{R}$] with a Polish group topology.
@@
Let $(Y,M)$ be a $\sigma$-finite measured space. We endow its measure algebra with the lower $M$-topology which is defined by taking as neighborhoods of an element $A\in\mathrm{MAlg}(Y,M)$ all sets of the form
@@
Note that the only neighborhood of the emptyset is the whole measure algebra, while $Y$ is contained in every neighborhood of every point. Moreover, $\{\emptyset\}$ is closed, and more generally $M$ is lower semi-continuous for the lower $M$-topology.
@@
Let $\epsilon>0$. Note that $(A_{1}\cap A_{2})\setminus(B_{1}\cap B_{2})\subseteq(A_{1}\setminus B_{1})\cup(A_{2}\setminus B_{2})$, so if if $M(A_{1}\setminus B_{1})<\epsilon$ and $M(A_{2}\setminus B_{2})<\epsilon$, then $M((A_{1}\cap A_{2})\setminus(B_{1}\cap B_{2}))<2\epsilon$ as wanted. ∎
@@
Let $(A_{n})_{n\in\mathbb{N}}$ be a family of elements of the measure algebra, then since $m$ is finite we may find $N$ such that $\mu((\bigcup_{n}A_{n})\setminus\bigcup_{n<N}A_{n}))<\epsilon$. Now if we take $B_{0},...,B_{N-1}$ such that $m(A_{i}\setminus B_{i})<\epsilon/N$, and any $B_{N},B_{N+1},$..., we see that $m((\bigcup_{n}A_{n})\setminus(\bigcup_{n}B_{n}))<2\epsilon$ as wanted. ∎
@@
Note that the countable union map is not continuous if we put the usual topology everywhere (indeed $X$ can easily be obtained as a limit of unions of sequences of sets which converge coordinatewise to $\emptyset$). Here is nevertheless a positive result for the usual topology which will be useful later on.